# Strong asymptotic giant branch stars’ spectral features in distant quiescent galaxies: Impact on galaxy evolution

### v1

@@ 2.3 The spectra of galaxies 8595 and 9025
The spectra of the other two fainter quiescent galaxies, 8595 (mag=24 at 3.6m, observed with the PRISM) and 9025 (mag=23 at 3.6m, observed with the medium-resolution gratings but rebinned to the PRISM resolution), have substantially lower SNR of 15 and 7. Nevertheless, the SNR is still sufficient for the direct detection of individual absorption features in object 8595 (Fig.4), some of the strongest also observed in object 9025 (Fig.5). This includes some TiO absorptions and CN spectral edges, and tentatively a C absorption.
Table 2 presents the results of the best spectral fits of objects 8595 and 9025, based on the different models. For object 8595, we carefully account for a neighbouring bright galaxy when measuring its photometry to estimate spectral aperture corrections, following a similar approach to what is done for object D36123 (see Methods). Again, M13 yields the best-fit spectral fits for both galaxies, with 0.8 for object 8595, and 1.7 for object 9025. The derived stellar masses differ at most by dex between the M13 and BC03 models, but the best-fit ages are even more strongly dependent on the specific model used (stellar age differences from 0.1 to 1 Gyr). In Table 2, the values denote the difference with respect to the M13 model. The small cumulative probability suggests the preference for the M13 model is statistically significant at the ( confidence level) in the overall spectral fit for these two galaxies.
@@ 3 Discussion
Overall, we conclude that the spectra of distant quiescent galaxies in this study show clear evidence for a substantial contribution of TP-AGB stars of both C and M type, at variance with what predicted by TP-AGB-poor and -light models. The models with significant TP-AGB contribution, such as M13 or M05, are in better agreement with the NIR features in the spectra of these high-redshift quiescent galaxies, and TP-AGB-mild models (i.e., M13) can fit their entire spectra better. These support our conclusion that the TP-AGB is the primary model ingredient driving the fitting of our spectrum.
@@
Previous work[18] based on high-quality NIR spectra of local spiral/Seyfert galaxies reached similar conclusions, favouring M05 models over BC03 models because of the TP-AGB component. Those local spectra cover a wide wavelength range similar to our high-redshift data and display a variety of features, several of which can only be attributed to TP-AGB stars. Previously, a detection of the CN 1.1m band in the NIR spectra of local AGNs was interpreted as a signature of intermediate age populations evolving along the TP-AGB phase [32]. Although these local observations are remarkable, JWST spectroscopy can directly observe massive, high-redshift galaxies, with low spread in stellar generations, for their stars to evolve to the TP-AGB phase in sync at ages 1 Gyr when the TP-AGB energetic contribution is maximum [3, 5]. This maximizes the contribution of the short TP-AGB phase (3 Myr for an individual star) to the integrated galaxy SED, with respect to systems with ongoing star formation and a lower mass fraction of stars in the appropriate age range. The dominance of the TP-AGB phase also implies that the NIR spectrum will display all its complexity and thus the SED fit becomes challenging.
@@
Indeed, we find that even for the best performing M13 model the is large for the high SNR spectrum of D36123 that shows stronger absorption features than the model. By separating optical (m) and NIR (m), we confirmed that it is the fit in the NIR spectrum leading to high values of , since we find a minimum reduced –11 in the optical spectrum for the four different models (C09 performing best in this case), while all reduced values remain large (40–60) when fitting only the NIR spectrum (see Table 1). More in detail, it appears that the rest-frame m range, where the strongest TiO features are present, is the most problematic region (see Table 1 and Fig.11 in the Methods) since when we exclude this range, both optical and NIR can be fitted to better than 2% accuracy by the M13 model (Fig.11; but not as well by BC03 and CO09).
Several model ingradients could be responsible for this discrepancy. First of all, M13 and M05 use empirical spectra for just the TP-AGB phase. For the other phase that strongly contribute to the NIR, the RGB, classical Kurucz-type theoretical spectra from model atmospheres are adopted, which - as it is well known - do not adequately match all NIR absorption features [39, 40]. Similarly, models based on the empirical RGB spectra perform better, but not for all features [39]. Notably, studies by different researchers show that all CO NIR spectral features from H to K in local massive early-type galaxies, whose NIR spectrum is dominated by RGB stars, are systematically stronger than those in the models based on empirical libraries [41]. This is also supported by the analysis of observed NIR spectra from local spirals and star-forming galaxies [42]. All this suggests that D36123 contain also stars that are more metal rich than those included in the model.
@@
Our model comparison suggests that extensions to the best-fitting models are needed to improve the match to massive galaxy spectra. The M13 models adopt local samples of C-rich and O-rich spectra, which may not cover the range of metallicity and abundance ratios hosted in distant massive ellipticals. The metallicity dependence in M13 models is included using a theoretical prescription [12], by which more metal-rich models have more fuel consumption in O-rich rather than in C-rich phases, but still using the same observed spectra. In particular, M13 models do not treat the relative contributions of C-rich and O-rich TP-AGB stars as a free parameter. Moreover, the calibration of stellar population models requires samples of stellar generations that are as simple as possible, i.e., star clusters, with available independent age and chemical composition from Color Magnitude Diagram (CMD) fitting and stellar spectroscopy. For the TP-AGB phase these calibrating clusters also need to span a wide age range, extending down to Myr. Star clusters with these properties exist in the Magellanic Clouds and have traditionally served to calibrate stellar population models [2, 43, 10]. The limitation is that these clusters have sub-solar metallicity and possibly solar or sub-solar element ratios.
@@
Another issue is a galaxy specific chemical pattern. It is well-known that massive galaxies, both locally [35, 44] as well as at high redshift [29, 45, 46], feature [/Fe]-enhanced stellar populations. The same galaxies are also found to host enhanced ratios of [C/Fe] and [N/Fe] [47, 48] and possibly also in Na [49, 46, 50]. Empirical libraries based on local, solar-scaled RGB spectra, and the population models based on them, may not reflect these chemical patterns, causing discrepancies. This would mirror a long-standing discrepancy between optical spectral features of massive galaxies and early population models [51], which was solved by -enhanced population models [52]. Galaxy spectra like those we present here, with young ages and super-solar metallicity, and possibly various non-solar element enhancements, are unique, with no known counterpart in the local universe. They challenge the models, but also offer the invaluable opportunity to calibrate them, with far-reaching benefit also for stellar evolution and model atmosphere theories.
@@
Astronomers depend on evolutionary population synthesis models to infer the timeline of galaxy formation and evolution. In turn, the model age scale depends on the input stellar evolution in terms of stellar tracks (and their input) and the energetics of stellar phases. We derive systematically younger ages and lower masses when fitting these spectra with TP-AGB-mild models (i.e., M13) versus TP-AGB-poor/-light models (i.e., BC03/C09), especially at high redshift. The age differences range from 100 Myr to 1 Gyr (see Tables 1 and 2). Younger galaxy ages can have implications for the inferred physics of galaxy evolution, such as gas accretion and star formation quenching. Also galaxy simulations adopt evolutionary population synthesis models as input, hence our results have an impact on the predicted spectra and photometry of simulated galaxies found in cosmological and other simulations [6]. In that case, the age is fixed within the simulation, but the simulated galaxy SEDs and the calibrations through comparisons to observations are obviously affected. The AGB stellar phase has also a major impact on galaxies dust production and chemical evolution particularly for heavy elements [53]. A higher fuel for this phase, as supported by our findings, has implications on guiding the relevant calculations for these processes [54], thus potentially having an even broader impact on galaxy evolution.
@@
The spectra presented here demonstrate the presence of TP-AGB spectral features, including those from carbon stars, in the integrated NIR light from high-redshift, quiescent galaxies. These features are first seen in their richness in galaxy spectra. Clearly, our very small galaxy sample does not allow us to infer the statistical impact of the TP-AGB emission on the derived stellar population properties for young quiescent galaxies in general. Previous photometric studies [5, 19] have shown that, on average, fitting distant galaxies using models with heavier TP-AGB contributions yields younger ages and lower stellar masses than TP-AGB-poor models, but also with significant galaxy-by-galaxy differences. The usual degeneracies between age, reddening and metallicity still apply to some extent, and for individual galaxies, the SED fitting results can depend on unknown details of the past star formation histories (for example). From our small sample, it is also difficult to evaluate how frequently these TP-AGB features are present in early galaxy spectra. NIRSpec observations for a larger, homogeneously selected sample of galaxies can address these questions and help us to quantify the broader consequences of TP-AGB emission for determining galaxy physical parameters. Such observations may also help guide further improvement of stellar population models.
@@
The 1st column denotes the rest-frame range of wavelength adopted in the spectral fit. The 3rd to 6th columns list the best-fitting physical properties derived by four models with different TP-AGB treatments. Specifically, the total stellar masses are based on the Chabrier[55] IMF. The best-fitting parameters and errors are derived from the distribution of , with the former obtained at -minimisation, and the errors determined after rescaling to . SFR/SFR is a measure of quiescence, where the SFR is derived from the best fit by adopting delay- models, and the SFR is corresponding to the SFR maximum at .
@@
For both objects, the spectral fit utilizes the entire rest-frame wavelength range. Additionally, to better constrain the age-sensitive wavelength around 4000Å, two additional HST bands (F606W+F814W) are employed simultaneously in the fit of object 9025 (see Methods). Measurements and errors are as in Tab.1, except that we did not rescale the reduced to 1. is computed relative to the M13 model.
4 Methods
@@ 4.1 NIRSpec observation and reduction
The NIRSpec data processing followed the same methodology as used for other CEERS NIRSpec observations [27, 28, 57, 58, 59]. The NIRSpec data reduction was based on the STScI Calibration Pipeline [60] version 1.8.5 and the Calibration Reference Data System (CRDS) mapping 1041 for objects 8595 and 9025, and CRDS mapping 1029 for D36123. There are three main stages in the reduction process. Briefly, the calwebb_detector1 pipeline module was utilized to correct for the detector 1/f noise, subtract the bias and dark current, generate the count-rate maps (CRMs) from the uncalibrated images. There is an improvement in the jump step to correct the cosmic ray “snowball” [28]. The stage two of the pipeline reduction is to create two-dimensional (2D) spectra with a rectified trace and flat slope from the generated CRMs. The background subtraction and slit loss correction of objects D36123 and 8595 were considered at this stage, while the slit loss correction (pathloss) was not taken into account for object 9025 as it is poorly centered in the shutter. We verified that better overall of its spectral fit can be obtained if adopting the empirical aperture corrections directly, which is the last step of the absolute flux calibrations (as discussed later in a dedicated subsection) to complement any incorrect calibration during the reduction process. At the third stage, we made use of optimised apertures for the extraction of the one-dimensional (1D) spectra in all cases. For object D36123, we adopted a 4-pixel extraction aperture, considering the effects of the Micro-Shutter Assembly [61] (MSA) bar shadows on some upper pixels in 2D spectra. We also considered 3- and 5-pixel apertures, with negligible impact on the results. The flux error for the extracted spectra was calculated by the JWST pipeline using an instrumental noise. We rescaled all flux errors multiplying by a factor of 1.5 to account for correlations induced by the pipeline [27].
@@ 4.3 Aperture correction
Spectra reduced with the NIRSpec pipeline include an aperture correction term that is derived assuming a point-source scenario. To account for the actual shapes of our galaxies, we integrate the observed spectra through imaging filter bandpasses to obtain synthetic photometry (i.e., ), and compare to the actual photometry (i.e., ) measured from imaging. We are implicitly assuming here that no strong color gradient biases the in-slit photometry with respect to the integrated galaxy photometry. This point has been verified by building PSF-matched color images of our targets and verifying that no color-gradients were apparent. We thus derive the aperture correction as . The top panel of Fig.8 shows the aperture correction for object D36123 derived in this way. We adopt a linear function to fit the trend with wavelength, as shown by the red line. There is a weak wavelength dependence (slope 0.02). The green line presents the rescaled aperture correction derived from the forward-modeling tool for NIRSpec Multi-Object Spectroscopic data, namely MSAFIT [67], which accounts for the complex geometry of the target galaxy, point spread function and pixelation of the NIRSpec instrument. This MSAFIT line aligns with the red linear relation. Regardless of the relation (linear vs. MSAFIT) or when adopting the extremes in the 1 range in the linear relation (the grey shaded region of D36123 in Fig.8) to correct the spectrum, the best-fit parameters and spectral fits show minimal changes. Hence we decided to finally rescale the observed spectrum of D36123 by adopting the linear relation. The same process is applied to the other two objects 8595 and 9025. Their spectra are rescaled multiplying by a constant value of 2.28 and a linear relation with a slope of -0.07, respectively.
@@ 4.4 Spectral resolution
The spectral resolution of the NIRSpec PRISM is relatively low. Based on the spectral sampling and for an unresolved source filling the slit, we can expect a resolution of 70 at 0.6m, declining to 30 at 1.1m, and rising again to 200 at 5m. For point/compact sources, the resolution of in-flight data could be 1.5-2.0 higher than that in the pre-launch approximation [67]. This would be helpful to identify some weak/narrower features, especially for galaxy D36123 where the SNR in the continuum is high. We adopt the MSAFIT [67] software to model the NIRSpec resolving power based on the actual galaxy spatial profile. The resolution modeled by MSAFIT software is consistent with that of a flat source in a slit width of 1.4 pixels (see green curve in the top-middle panel of Fig.9).
The intrinsic dispersion of our target early-type galaxies (ETGs) need to be accounted next. The velocity dispersion can be estimated using local stellar mass vs. velocity dispersion relations (i.e., -) relation [35]. We convert to the full width at half maximum (FWHM) and define (where FWHM). The intrinsic broadening of an ETG with a velocity dispersion 130km/s (a ballpark number for the typical masses in our study) corresponds to R1000 (see the grey line in the top-middle panel of Fig.9), quite negligible compared to the instrumental spread.
Finally, we estimated the template model’s resolution from their spectral sampling, adopting 2 spectral bin widths as their (Nyquist) sampling. The sampling (hence resolution) is changing in steps, higher in the optical than in the NIR. Redshifting will affect where these steps actually apply for the observed galaxies. The corresponding resolution in the observed frame for D36123 (assuming an average 1.077), is shown in the top-middle panel of Fig.9.
To match the resolution of the models with the observed spectrum, we need to account for all three impacting factors: prism dispersion (dominant term), the wavelength sampling of the spectral template models, and the velocity dispersion of the ETG. We add and subtract in quadrature the corresponding terms at fixed . The broadening needed to be applied to models to match the observed spectra is provided by the formula:
where refers to the model. The final smoothing functions applied to models are equivalent to resolutions among ranging from R 60 to 450 across observed wavelength range of 0.6–3.3m, as depicted in the top right panel of Fig.9. Beyond 3.3m we are limited by the models resolution and therefore no further smoothing is applied to them. We repeat similar procedures to obtain the appropriate smoothing kernels for objects 8595 and 9025, shown in the middle and bottom panels of Fig.9. For both, the prism resolution is based on 1.4-pixel elements, and the spectra in different models was resampled in every 2 spectral bin widths, similar to what is done for object D36123.

### v2

@@ 2.3 The spectra of galaxies 8595 and 9025
The spectra of the other two fainter quiescent galaxies, 8595 (magAB=24 at 3.6m, observed with the PRISM) and 9025 (magAB=23 at 3.6m, observed with the medium-resolution gratings but rebinned to the PRISM resolution), have substantially lower SNR, 15 and 7. Nevertheless, the SNR is still sufficient to detect individual absorption features in object 8595 (Fig.4a). Some of the strongest features are also observed in object 9025 (Fig.4b), including TiO absorption lines, CN spectral edges, and tentatively a C2 absorption.
Table 2 presents the results of the spectral fitting of objects 8595 and 9025. For object 8595, we carefully account for a neighbouring bright galaxy when measuring its photometry to estimate spectral aperture corrections, following a similar approach to what is done for object D36123 (see Methods). Again, M13 yields the best-fit spectral fits for both galaxies, with 0.9 for object 8595, and 1.7 for object 9025. The derived stellar masses differ at most by dex between the M13 and BC03 models, but the best-fit ages are even more strongly dependent on the specific model used (0.1-1 Gyr). In Table 2, the values denote the difference with respect to the M13 model. The small cumulative probability suggests the preference for the M13 model is statistically significant in the overall spectral fitting for these two galaxies.
@@ 3 Discussion
Overall, we conclude that the spectra of distant quiescent galaxies in this study show clear evidence of substantial contribution of both C- and M-type TP-AGB stars. The models with significant TP-AGB contribution, such as M13 or M05, are in better agreement with NIR features in the galaxies’ spectra, while TP-AGB-mild models (i.e., M13) can fit their full spectra better, supporting our conclusion that the TP-AGB is the primary model ingredient driving the quality of the fits.
@@
Previous work[18] based on NIR spectra of local spiral/Seyfert galaxies reached similar conclusions, favouring M05 models over BC03 models because of the TP-AGB component. Those local spectra cover a wide wavelength range similar to our high-redshift data and display various features, several of which can only be attributed to TP-AGB stars. Previously, detection of the CN 1.1m band in the NIR spectra of local AGNs was interpreted as a signature of intermediate age populations evolving along the TP-AGB phase[30]. Although these local observations are remarkable, JWST spectroscopy can observe massive galaxies at high redshift, with low spread in stellar generations for their stars to evolve through the TP-AGB phase in sync at age 1 Gyr, when the TP-AGB energetic contribution is maximum[4, 6]. This maximizes the contribution of the short TP-AGB phase (3 Myr for an individual star) to the integrated galaxy SED, with respect to systems with ongoing star formation and a lower mass fraction of stars in the appropriate age range.
@@
Indeed, we find that even for the best-performing M13 model the is large for the high SNR spectrum of D36123, which shows stronger absorption features than the model. By separating optical (0.5m) and NIR (0.5m), we confirmed that it is the fit in the NIR spectrum leading to high values of , since we find a minimum reduced 9-11 in the optical spectrum for the four different models (C09 performing best in this case), while the reduced remains large (40-60) when fitting only the NIR spectrum (Table 1). Interestingly, excluding from the fit the range 0.5-1.0m, which contains the strongest TiO features, the M13 and M05 models largely over-perform the others when reproducing solely this region, pointing to a more consistent treatment of this spectral range heavily affected by TP-AGB stars (Extended Data Fig.4).
@@
In any case, overall remains always large for any specific model or spectral range, suggesting that improvements to the models are necessary. Focusing on the best-fitting models, M13 (as M05) uses empirical spectra for just the TP-AGB phase. For the other phases contributing to the NIR, as the RGB, classical Kurucz-type theoretical spectra from model atmospheres are adopted, which - as it is well known - do not adequately match all NIR absorption features[37, 38]. On the other hand, models based on the empirical RGB spectra perform better, but not for all features[37], as we also show here. Notably, various studies show that all CO NIR spectral features from H to K in local massive early-type galaxies, whose NIR spectrum is dominated by RGB stars, are systematically stronger than those in models based on empirical libraries[39]. This is also observed in the NIR spectra of local spirals and star-forming galaxies[40]. This evidence suggests that D36123 contains also stars that are more metal-rich than those included in the model. Also, the M13 models adopted local samples of C-rich and O-rich spectra, which may not cover the range of metallicity and abundance ratios hosted in distant massive ellipticals. The metallicity dependence of the relative contributions of C-rich and O-rich TP-AGB stars in M13 models is included using a theoretical prescription[2], where more metal-rich models have more fuel consumption in O-rich rather than in C-rich phases, but still using the same observed spectra. Moreover, the calibration of stellar population models requires samples of stellar generations that are as simple as possible, i.e., star clusters, with available independent age and chemical composition from Color Magnitude Diagram fitting and stellar spectroscopy. For the TP-AGB phase, these calibrating clusters also need to span a wide age range, extending down to 100 Myr. Star clusters with these properties exist in the Magellanic Clouds and have traditionally served to calibrate stellar population models[3, 41, 11]. The limitation is that these clusters have sub-solar metallicity and possibly solar or sub-solar element ratios.
@@
Another issue is a galaxy-specific chemical pattern. It is well-known that massive galaxies, both locally[33] as well as at high redshift[27, 42], feature [/Fe]-enhanced stellar populations. The same galaxies are also found to host enhanced ratios of [C/Fe] and [N/Fe][43, 44] and possibly also in Na[45, 46]. Empirical libraries based on local, solar-scaled RGB spectra, and the population models based on them, may not reflect these chemical patterns, causing discrepancies. This would mirror a long-standing discrepancy between optical spectral features of massive galaxies and early population models[47], which was solved by [/Fe]-enhanced population models[48]. Galaxy spectra like those we present here, with young ages and super-solar metallicity, and possibly various non-solar element enhancements, are unique, with unknown counterpart in the Universe. They challenge the models but also offer the invaluable opportunity to calibrate them, with far-reaching benefit also for stellar evolution and model atmosphere theories.
@@
Inferring the timeline of galaxy evolution depends on population synthesis models. In turn, the model age scale depends on the input stellar evolution in terms of stellar tracks (and their input) and the energetics of stellar phases. We derive systematically younger ages and lower masses when fitting these spectra with TP-AGB-mild models (M13) versus TP-AGB-poor/-light models (BC03/C09), especially at high redshift. The age differences range from 100 Myr to 1 Gyr (see Tables 1 and 2). Younger galaxy ages can have implications for the inferred physics of galaxy evolution, such as gas accretion and star formation quenching. Also, galaxy simulations adopt evolutionary population synthesis models as input, hence our results impact the predicted spectra and photometry of simulated galaxies found in cosmological and other simulations[7]. In that case, the age is fixed within the simulation, but the simulated galaxy SEDs and the calibrations through comparisons to observations are obviously affected. The AGB stellar phase has also a major impact on galaxies’ dust production and chemical evolution particularly for heavy elements[49]. More fuel for this phase, as supported by our findings, has implications on the relevant calculations for these processes[50], thus potentially having an even broader impact on galaxy evolution.
@@
The spectra presented here demonstrate the presence of TP-AGB spectral features, including those from carbon stars, in the integrated NIR light of high-redshift, quiescent galaxies. For the first time, these features are observed in all their complexity in galaxy spectra. Clearly, our very small galaxy sample does not allow us to infer, statistically, the impact of the TP-AGB emission on the derived stellar population properties of young quiescent galaxies. Previous photometric studies[6, 19] have shown that, on average, fitting distant galaxies using models with heavier TP-AGB contributions yields younger ages and lower stellar masses than TP-AGB-poor models, but with significant galaxy-by-galaxy differences. The usual degeneracies between age, reddening and metallicity still apply to some extent and, for individual galaxies, the SED fitting results can depend on unknown details of the past SFHs (for example). From our small sample, evaluating how frequently these TP-AGB features are present in early galaxy spectra is also difficult. NIRSpec observations of a larger, homogeneously selected sample of galaxies as luminous as D36123 (which are common, even over a single NIRSpec pointing) can address these questions and help us quantify the broader consequences of TP-AGB emission in determining galaxies’ physical parameters. Such observations may also help guide further improvement of stellar population models.
@@
| 446.7 | 308.4 | 157.9 | 172.1 | ||
The 1st column denotes the rest-frame range of wavelength adopted in the spectral fit. The 3rd to 6th columns list the best-fitting physical properties derived by four models with different TP-AGB treatments. Specifically, the total stellar masses are based on the Chabrier[51] IMF. Age means the mass-weighted age, i.e., an average look-back time weighted by the SFH. The best-fitting parameters and errors are derived from the distribution of , with the former obtained at -minimisation, and the errors determined after rescaling to . SFRbest/SFRpeak is a measure of quiescence, where the SFRbest is derived from the best fit by adopting delay- models, and the SFRpeak is corresponding to the SFR maximum at . presents the reduced within the masked 0.5-1m region during the spectral fit.
@@
For both objects, the spectral fit utilizes the entire rest-frame wavelength range. Additionally, to better constrain the age-sensitive wavelength around 4000Å, two additional HST bands (F606W+F814W) are employed simultaneously in the fit of object 9025 (see Methods). Measurements and errors are as in Table 1, except that we did not rescale the reduced to 1. is computed relative to the M13 model.
Methods
@@ 3.1 NIRSpec observation and reduction
The NIRSpec data processing followed the same methodology as used for other CEERS NIRSpec observations[21, 22, 53, 54, 55]. The NIRSpec data reduction was based on the STScI Calibration Pipeline[56] version 1.8.5 and the Calibration Reference Data System (CRDS) mapping 1041 for objects 8595 and 9025, and CRDS mapping 1029 for D36123. There are three main stages in the reduction process. Briefly, the calwebb_detector1 pipeline module was utilized to correct for the detector 1/f noise, subtract the bias and dark current, and generate the count-rate maps (CRMs) from the uncalibrated images. There is an improvement in the jump step to correct the cosmic ray “snowball”[22]. The stage two of the pipeline reduction is to create two-dimensional (2D) spectra with a rectified trace and flat slope from the generated CRMs. The background subtraction and slit loss correction of objects D36123 and 8595 were considered at this stage, while the slit loss correction (pathloss) was not taken into account for object 9025 as it is poorly centered in the shutter. We verified that better overall of its spectral fit can be obtained if adopting the empirical aperture corrections directly, which is the last step of the absolute flux calibrations (as discussed later in a dedicated subsection) to complement any incorrect calibration during the reduction process. At the third stage, we made use of optimised apertures for the extraction of the one-dimensional (1D) spectra in all cases. For object D36123, we adopted a 4-pixel extraction aperture, considering the effects of the Micro-Shutter Assembly[57] (MSA) bar shadows on some upper pixels in 2D spectra. We also considered 3- and 5-pixel apertures, with negligible impact on the results. The flux error for the extracted spectra was calculated by the JWST pipeline using an instrumental noise. We rescaled all flux errors multiplying by a factor of 1.5 to account for correlations induced by the pipeline[21].
@@ 3.3 Aperture correction
Spectra reduced with the NIRSpec pipeline include an aperture correction term that is derived assuming a point-source scenario. To account for the actual shapes of our galaxies, we integrate the observed spectra through imaging filter bandpasses to obtain synthetic photometry (i.e., ), and compare to the actual photometry (i.e., ) measured from imaging. We are implicitly assuming here that no strong color gradient biases the in-slit photometry with respect to the integrated galaxy photometry. This point has been verified by building PSF-matched color images of our targets and verifying that no color gradients were apparent. We thus derive the aperture correction as . The top panel of Extended Data Fig.2 shows the aperture correction for object D36123 derived in this way. We adopt a linear function to fit the trend with wavelength, as shown by the red line. There is a weak wavelength dependence (slope 0.02). The green line presents the rescaled aperture correction derived from the forward-modeling tool for NIRSpec Multi-Object Spectroscopic data, namely MSAFIT[63], which accounts for the complex geometry of the target galaxy, point spread function and pixelation of the NIRSpec instrument. This MSAFIT line aligns with the red linear relation. Regardless of the relation (linear vs. MSAFIT) or when adopting the extremes in the 1 range in the linear relation (the grey shaded region of D36123 in Extended Data Fig.2) to correct the spectrum, the best-fit parameters and spectral fits show minimal changes. Hence we decided to finally rescale the observed spectrum of D36123 by adopting the linear relation. The same process is applied to the other two objects 8595 and 9025. Their spectra are rescaled multiplying by a constant value of 2.28 and a linear relation with a slope of -0.07, respectively.
@@ 3.4 Spectral resolution
The spectral resolution of the NIRSpec PRISM is relatively low. Based on the spectral sampling and for an unresolved source filling the slit, we can expect a resolution of 70 at 0.6m, declining to 30 at 1.1m, and rising again to 200 at 5m. For point/compact sources, the resolution of in-flight data could be 1.5-2.0 higher than that in the pre-launch approximation[63]. This would be helpful to identify some weak/narrower features, especially for galaxy D36123 where the SNR in the continuum is high. We adopt the MSAFIT[63] software to model the NIRSpec resolving power based on the actual galaxy spatial profile. The resolution modelled by MSAFIT software is consistent with that of a flat source in a slit width of 1.4 pixels (see green curve in the top-middle panel of Extended Data Fig.3).
The intrinsic dispersion of our target early-type galaxies (ETGs) needs to be accounted next. The velocity dispersion can be estimated using local stellar mass vs. velocity dispersion relations (i.e., -) relation[33]. We convert to the full width at half maximum (FWHM) and define (where FWHM). The intrinsic broadening of an ETG with a velocity dispersion 130km/s (a ballpark number for the typical masses in our study) corresponds to R1000 (see the grey line in the top-middle panel of Extended Data Fig.3), quite negligible compared to the instrumental spread.
Finally, we estimated the template model’s resolution from their spectral sampling, adopting 2 spectral bin widths as their (Nyquist) sampling. The sampling (hence resolution) is changing in steps, higher in the optical than in the NIR. Redshifting will affect where these steps actually apply for the observed galaxies. The corresponding resolution in the observed frame for D36123 (using 1.08), is shown in the top-middle panel of Extended Data Fig.3.
To match the resolution of the models with the observed spectrum, we need to account for all three impacting factors: prism dispersion (dominant term), the wavelength sampling of the spectral template models, and the velocity dispersion of the ETG. We add and subtract in quadrature the corresponding terms at fixed . The broadening needed to be applied to models to match the observed spectra is provided by the equation:
where refers to the model. The final smoothing functions applied to models are equivalent to resolutions ranging from R 50 to 400 across the observed wavelength range of 0.6–3.3m, as depicted in the top right panel of Extended Data Fig.3. Beyond 3m for some models we are limited by the models’ resolution. No further smoothing is applied to those. We repeat a similar procedures to obtain the appropriate smoothing kernels for objects 8595 and 9025, shown in the middle and bottom panels of Extended Data Fig.3. For both, the prism resolution is based on 1.4-pixel elements, and the spectra in different models were resampled in every 2 spectral bin widths, similar to what is done for object D36123.